OpenStudy (pinknabastak):

ASAP!!!!! HELLLPPPPPPP!!!!! please! need the answer ASAP find the 8th term:

OpenStudy (pinknabastak):

\[(2x-4)^{13}\]

OpenStudy (pinknabastak):

@ganeshie8 hey can you please help me get the answer

OpenStudy (pinknabastak):

@BAdhi HELPPP!

OpenStudy (badhi):

by 8th term I guess you are asking the coefficient of the \(x^8\). Use binomial expansion \((a+b)^n =\displaystyle \sum\limits_{r=0}^n\binom {n}{r}a^rb^{n-r} \)

OpenStudy (pinknabastak):

uhh i dont understand that

OpenStudy (badhi):

Haven't you learned about Binomial theorem? http://en.wikipedia.org/wiki/Binomial_theorem

OpenStudy (nirmalnema):

you have to find the 8th term of what

OpenStudy (pinknabastak):

every time someone tries to explain this to me i just dont get it I've been trying to get the answer to this problem for the past 2 days

OpenStudy (pinknabastak):

the 8th term of that equation

OpenStudy (badhi):

First, there is no such thing as 8th term in this expression and second, there isn't any equation here. @pinknabastak, i think you want to expand this expression and find the coefficient of the \(x^8\) term am I right?

OpenStudy (pinknabastak):

i guess my teacher didnt explain at all it just says find the 8th term... but i think thats what he wants

OpenStudy (***[isuru]***):

hey is this in a book or is it an online one ? either way if u can post a snapshot of the original question it would be lot easy to understand what is missing here

OpenStudy (pinknabastak):

its a worksheet

OpenStudy (pinknabastak):

acually my final test i just didnt take it last year and my teacher gave it to me and told me to finish it this break and hell give me credit for the class for the whole year so its kindaa important to finish it

OpenStudy (pinknabastak):

sorry the pic is backwards

OpenStudy (***[isuru]***):

no prob... try this site... http://www.ltcconline.net/greenl/courses/154/seqser/binomial.htm

OpenStudy (***[isuru]***):

follow it and u will able to find the 8th term

OpenStudy (pinknabastak):

huh?

OpenStudy (nirmalnema):

just wait i m trying

OpenStudy (***[isuru]***):

do u know how to expand an expression like \[(x + 2)^{2} or ( y + 6)^{3} or (a+x)^{n}\] ?

OpenStudy (pinknabastak):

umm (x+2)(x+2) or (y+6)(y+6)(y+6)

OpenStudy (***[isuru]***):

that;s the easy way... but do u know how to do it like this (x +y)^2 = x^2 + 2xy + y^2 ?

OpenStudy (nirmalnema):

\[(2x-4)^{13}= [(2x)^{13}]+[13 C 1 *(2x)^{12} *(-4)]+[(2x)^{11}*13C2*(-4)^{2}]---------and so on\]

OpenStudy (pinknabastak):

uh no...is that FOIL?

OpenStudy (pinknabastak):

@nirmalnema what do i do with that?

OpenStudy (nirmalnema):

this is the way of expanding any term in binomial here nCr = n! / r!(n-r)! its the combination of terms./. which i have used.. and you have to do it futher to get the 8 th term

OpenStudy (pinknabastak):

how do i do that?

OpenStudy (nirmalnema):

u have to use the formula of binomial expansion

OpenStudy (pinknabastak):

honestly i have no idea how to use that..can you just tell me the 8th term..cuz obviously you know it and i obviously i cant figure it out and i will MORE then likely never use this again....

OpenStudy (nirmalnema):

according to this the answer will be \[^{13} C _{7} *(2x)^{6}*(-4)^{7}\] is the 8th term

OpenStudy (pinknabastak):

do i put the C down to...?

OpenStudy (nirmalnema):

what i dont understand what u said

OpenStudy (pinknabastak):

do i write down that whole thing....like whats the "C" for? in the begining of the answer

OpenStudy (anonymous):

What grade math is that?

OpenStudy (pinknabastak):

junior/senior class....its Algebra 2

OpenStudy (anonymous):

Oh wow, im taking Geometry 9th grade. Sorry I can't help lol

OpenStudy (nirmalnema):

yup you can write it like that also..if you need i can tell you in the simplified form

OpenStudy (pinknabastak):

yes please! :)

OpenStudy (pinknabastak):

@Sally♡ haha yea next year for you

OpenStudy (nirmalnema):

1716 * 64x^7 * (-16384)

OpenStudy (pinknabastak):

THANK YOU SOOOO MUCH!!!

OpenStudy (nirmalnema):

your welcome... :) ....!!!!!!!!!!